OpenStudy (anonymous):

optimization calculus problem help. Show that among all rectangles with 8-m perimeter, the one that is largest is a square.

OpenStudy (anonymous):

 let perimeter = 2x + 2y = 8 maximize perimeter (take derivative of it) set it = 0

OpenStudy (anonymous):

the derivative of it would be 2+2=0

OpenStudy (anonymous):

do you mean the one that is the largest area?

OpenStudy (anonymous):

yeah largest area

OpenStudy (anonymous):

if thats what you mean you have to maximize the area equation A = xy

OpenStudy (anonymous):

where 2x + 2y = 8, so 2y = 8 - 2x y = 4 - x A = x*y A = x (4 - x) maximize this one

OpenStudy (anonymous):

4-2x=0

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

so x = 2 that means the perimeter is p = 2x + 2y = 8 p = 4 + 2y = 8 solve for y

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

perfect!

OpenStudy (anonymous):

how would i write the answer?

OpenStudy (anonymous):

well the work you showed proved that the value of x when the area is maximized is x = 2 then you solved for y using x = 2, and found y = 2 the dimensions of the rectangle is x by y, or 2 by 2, which is a square that means to maximize the area, the rectangle has to be a square

OpenStudy (anonymous):

thank you so much for your help!